# De Finetti’s problem with fixed transaction costs and regime switching

### v1

@@ Appendix A Proofs
$\zeta(z^{\prime}_{1},z^{\prime}_{2})=1/g^{\prime}(z^{\prime}_{2})>1/g^{\prime}(z_{2})=\zeta(z_{1},z_{2}),$
which means that any $(z_{1},z_{2})\in\mathcal{N}$ such that $z_{2}\in(a,x_{0}]$ satisfies $(z_{1},z_{2})\notin\mathcal{M}_{\zeta}$. Hence, (1) holds true. The other claims can be proved by similar arguments combined with the definition of $x_{1}$ and $x_{2}$. We hence omit their proofs. By the above claims (1)-(6), we know that $\mathcal{M}_{\zeta}\subseteq\cup_{i=1}^{4}\mathcal{R}_{i}$, where
@@
$\mathcal{R}_{1}$ =$=\begin{cases}\{(\tilde{z}_{1},\tilde{z}_{2})\},\,\text{ if }\beta\leq\omega_{1}(x_{2}),\\ \emptyset,\text{ if else},\end{cases}$
and
$\cup_{i=2}^{4}\mathcal{R}_{i}$ =$=\begin{cases}\{(\bar{z}_{1},\bar{z}_{2})\},\,\text{ if }\beta<\omega_{2}(x_{1}),\\ \{(\hat{z}_{1}:=(g^{\prime})^{-1}_{1}(g^{\prime}(\omega_{2}^{-1}(\beta))),\hat{z}_{2}:=\omega_{2}^{-1}(\beta))\},\,\text{ if }\beta>\omega_{2}(x_{1}),\\ \{(\bar{z}_{1},\bar{z}_{2}),(\hat{z}_{1},\hat{z}_{2})\},\,\text{ if }\beta=\omega_{2}(x_{1}).\end{cases}$
Suppose $\beta<\omega_{2}(x_{1})$.
  • If $g^{\prime}(\omega_{1}^{-1}(\beta))<g^{\prime}(\omega_{2}^{-1}(\beta))$, then $\emptyset\neq\mathcal{M}_{\zeta}\subseteq\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2})\}$, and
@@
  • If $g^{\prime}(\omega_{1}^{-1}(\beta))>g^{\prime}(\omega_{2}^{-1}(\beta))$, then $\emptyset\neq\mathcal{M}_{\zeta}\subseteq\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2})\}$, and
@@
  • If $g^{\prime}(\omega_{1}^{-1}(\beta))=g^{\prime}(\omega_{2}^{-1}(\beta))$, then $\emptyset\neq\mathcal{M}_{\zeta}\subseteq\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2})\}$, and
@@
Suppose $\beta=\omega_{2}(x_{1})$.
  • Note $g^{\prime}(\omega_{1}^{-1}(\beta))>g^{\prime}(\omega_{2}^{-1}(\beta))$ cannot hold in this case.
  • If $g^{\prime}(\omega_{1}^{-1}(\beta))=g^{\prime}(\omega_{2}^{-1}(\beta))$, then $x_{2}\leq x_{1}$, $\emptyset\neq\mathcal{M}_{\zeta}\subseteq\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2}),(\hat{z}_{1},\hat{z}_{2})\}$, $\hat{z}_{1}=\tilde{z}_{1}$ and $\hat{z}_{2}=\tilde{z}_{2}$ (since $\omega_{1}\equiv\omega_{2}$ on $[x_{1}\vee x_{2},\infty)$), and
@@
  • If $g^{\prime}(\omega_{1}^{-1}(\beta))<g^{\prime}(\omega_{2}^{-1}(\beta))$, then $x_{1}<x_{2}$, $\emptyset\neq\mathcal{M}_{\zeta}\subseteq\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2}),(\hat{z}_{1},\hat{z}_{2})\}$, and
@@
Suppose $\beta>\omega_{2}(x_{1})$.
  • If either of the following cases holds
@@
      $x_{1}<x_{2}$ and $\hat{z}_{2}\geq x_{2}$,
    then $g^{\prime}(\omega_{1}^{-1}(\beta))=g^{\prime}(\omega_{2}^{-1}(\beta))$ (since $\omega_{1}\equiv\omega_{2}$ on $[x_{1}\vee x_{2},\infty)$), and consequently $\hat{z}_{1}=\tilde{z}_{1}$, $\hat{z}_{2}=\tilde{z}_{2},$ and $\cup_{i=1}^{4}\mathcal{R}_{i}=\{(\tilde{z}_{1},\tilde{z}_{2})\}$. Hence, $\mathcal{M}_{\zeta}=\{(\tilde{z}_{1},\tilde{z}_{2})\}$.
  • If $x_{1}<x_{2}$ and $\hat{z}_{2}<x_{2}$, then
@@
$=[g^{-}(0)-g^{+}(0)](\theta_{1}^{+})^{2}>0.$
It is seen that the function $g^{\prime\prime}(x)$ is strictly increasing on $(a,\infty)$, which together with above implies that $g^{\prime\prime}(x)>0$ on $(a,\infty)$. The proof is complete.
@@
      If $a>a_{1}$, one has $g^{\prime\prime}(x)<0$ on $[0,a_{1})$ and $g^{\prime\prime}(x)>0$ on $(a_{1},a]$.
@@
    $=[g^{-}(0)-g^{+}(0)](\theta_{1}^{+})^{2}>0,$
    where we have used the facts that $\theta_{1}^{+}\leq\theta_{2}^{+}$ and $g^{-}(0)>g^{+}(0)>0$. It is seen that the function $g^{\prime\prime}(x)$ is strictly increasing on $(a,\infty)$, which together with above implies that $g^{\prime\prime}(x)>0$ on $(a,\infty)$.
@@
We exclusively present the proof for case (iv) of Subsection 3.1, as proofs for the remaining cases of Subsections 3.1-3.4 adhere to patterns that are either similar or simpler in nature. Recall that $\omega_{1}$ coincides with $\omega_{2}$ on $[x_{1}\vee x_{2},\infty)$. By Theorem 3.1, one has
@@
    $z_{2}-z_{1}=\begin{cases}\omega_{2}^{-1}(\beta)-(g^{\prime})_{3}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),&\beta\in[\omega_{1}(x_{2}),\omega_{2}(x_{1})),\\ \omega_{2}^{-1}(\beta)-(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),&\beta\in[\omega_{2}(x_{1}),\infty),\\ \omega_{1}^{-1}(\beta)-(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{1}^{-1}(\beta))),&\beta\in[0,\omega_{1}((g^{\prime})_{2}^{-1}(g^{\prime}(x_{0})))),\end{cases}$
    which implies that $z_{2}-z_{1}$ is increasing in $\beta$ on $[0,\omega_{1}((g^{\prime})_{2}^{-1}(g^{\prime}(x_{0}))))\cup[\omega_{2}(x_{2}),\infty)$.
@@
    $z_{2}-z_{1}=\omega_{1}^{-1}(\beta)-(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{1}^{-1}(\beta))),\quad\beta\in[0,\omega_{1}((g^{\prime})_{2}^{-1}(g^{\prime}(x_{0}))))\cup[\omega_{1}(x_{1}),\infty),$
    which implies that $z_{2}-z_{1}$ is increasing in $\beta$ on $[0,\omega_{1}((g^{\prime})_{2}^{-1}(g^{\prime}(x_{0}))))\cup[\omega_{1}(x_{1}),\infty)$.
@@
    $z_{2}-z_{1}=\begin{cases}\omega_{2}^{-1}(\beta)-(g^{\prime})_{3}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),&\beta\in[0,\omega_{2}(a_{5})]\cup[\omega_{2}(x_{2}),\omega_{2}(x_{1})),\\ \omega_{2}^{-1}(\beta)-(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),&\beta\in[\omega_{2}(x_{1}),\infty),\end{cases}$
    which implies that $z_{2}-z_{1}$ is increasing in $\beta$ on $[0,\omega_{2}(a_{5})\cup[\omega_{2}(x_{2}),\infty)$.
@@
    $z_{2}-z_{1}=\begin{cases}\omega_{2}^{-1}(\beta)-(g^{\prime})_{3}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),&\beta\in[0,\omega_{2}(a_{5})],\\ \omega_{1}^{-1}(\beta)-(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{1}^{-1}(\beta))),&\beta\in[\omega_{2}(x_{1}),\infty),\end{cases}$
    which implies that $z_{2}-z_{1}$ is increasing in $\beta$ on $[0,\omega_{2}(a_{5})]\cup[\omega_{2}(x_{1}),\infty)$.
In the sequel, we only check the increasing property of $\max_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ and $\min_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ (with respect to $\beta$) over the interval $[\omega_{1}((g^{\prime})_{2}^{-1}(g^{\prime}(x_{0}))),\omega_{1}(x_{2})]$ under the above Case (a), because the corresponding proofs needed for the above Cases (b)-(d) are quite similar.
@@
Hence, it holds that $[\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),\omega_{1}(x_{2})]=\Lambda^{+}\cup\Lambda^{-}\cup\Lambda^{0}.$ We next demonstrate that $\Lambda^{\pm}$ can each be expressed as a countable union of disjoint open intervals, and, $\Lambda^{0}$ can be expressed as a countable union of disjoint closed intervals. For any $\beta_{0}\in\Lambda^{+}$, we have $g^{\prime}(\omega^{-1}_{1}(\beta_{0}))>g^{\prime}(\omega^{-1}_{2}(\beta_{0}))$. According to the sign-preserving property of continuous functions, there exists a constant $\epsilon>0$ such that $(\beta_{0}-\epsilon,\beta_{0}+\epsilon)\subseteq[\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),\omega_{2}(x_{2})]$ and for any $\beta\in(\beta_{0}-\epsilon,\beta_{0}+\epsilon)$, it holds that $g^{\prime}(\omega^{-1}_{1}(\beta))>g^{\prime}(\omega^{-1}_{2}(\beta))$. Denote
@@
[…]
@@
[…]
With the continuity of $g^{\prime}$ on $\mathbb{R}_{+}$ and $(\omega^{-1}_{i})_{i=1,2}$ on $[\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),\omega_{1}(x_{2})]$, we derive that $g^{\prime}(\omega^{-1}_{1}(\beta_{0}-\underline{\epsilon}(\beta_{0})))=g^{\prime}(\omega^{-1}_{2}(\beta_{0}-\underline{\epsilon}(\beta_{0})))$ and $g^{\prime}(\omega^{-1}_{1}(\beta_{0}+\overline{\epsilon}(\beta_{0})))=g^{\prime}(\omega^{-1}_{2}(\beta_{0}+\overline{\epsilon}(\beta_{0})))$. Additionally, it holds that $\beta_{0}-\underline{\epsilon}(\beta_{0})\leq\beta_{0}-\epsilon<\beta_{0}<\beta_{0}+\epsilon\leq\beta_{0}+\overline{\epsilon}(\beta_{0}),$ and $(\beta_{0}-\underline{\epsilon}(\beta_{0}),\beta_{0}+\overline{\epsilon}(\beta_{0}))$ is a non-empty subset of $\Lambda^{+}$. In this vein, for any $\beta_{0}\in\Lambda^{+}$, one can find a non-empty open subset of $\Lambda^{+}$ that contains $\beta_{0}$. In addition, according to the denseness of rational numbers in the set of real numbers, there is at least one rational number in $(\beta_{0}-\underline{\epsilon}(\beta_{0}),\beta_{0}+\overline{\epsilon}(\beta_{0}))$. However, rational numbers are countable, implying that $\Lambda^{+}$ is a countable union of disjoint open intervals. Using similar arguments, one can obtain that $\Lambda^{-}$ is also a countable union of disjoint open intervals. Furthermore, it follows from $[\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),\omega_{1}(x_{2})]=\Lambda^{+}\cup\Lambda^{-}\cup\Lambda^{0}$ that $\Lambda^{0}$ is a countable union of disjoint closed intervals. More specifically, the three sets $\Lambda^{+}$, $\Lambda^{-}$ and $\Lambda^{0}$ can be expressed as
@@
We next demonstrate that $\max_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ and $\min_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ are indeed increasing functions of $\beta$ on $[\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),\omega_{1}(x_{2})]$ under Case (a). Actually, by definition, one knows that if $\beta=\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0})))$, then $g^{\prime}(\omega^{-1}_{1}(\beta))=g^{\prime}(x_{0})<g^{\prime}(\omega^{-1}_{2}(\beta))$, which yields $\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0})))\in\Lambda^{-}$, implying that $\Lambda^{-}\neq\emptyset$. Similarly, if $\beta=\omega_{1}(x_{2})$, then $g^{\prime}(\omega^{-1}_{1}(\beta))=g^{\prime}(x_{2})>g^{\prime}(\omega^{-1}_{2}(\beta))$, which yields $\omega_{1}(x_{2})\in\Lambda^{+}$, implying that $\Lambda^{+}\neq\emptyset.$ Then by (A.16), a point $(z_{1},z_{2})\in\mathcal{M}_{\zeta}$ satisfies that
@@
It can be checked that the functions $\varphi_{1}$ and $\varphi_{2}$ are strictly increasing on $[g^{\prime}(x_{0}),g^{\prime}(x_{2})]$, and inherit differentiability from $\omega_{1}$, $\omega_{2}$, $(g^{\prime})_{2}^{-1}$ and $(g^{\prime})_{4}^{-1}$. In addition, it holds that
@@
$\omega_{2}((g^{\prime})^{-1}_{4}(u))=\omega_{1}((g^{\prime})^{-1}_{2}(u)),\quad u\in[\varphi_{1}^{-1}(t^{-}_{j}),\varphi_{1}^{-1}(t^{+}_{j})].$
This implies that for any $\beta\in(t^{-}_{j},t^{+}_{j})$ and $v=\varphi_{1}^{-1}(\beta)=g^{\prime}(\omega^{-1}_{1}(\beta))$,
@@
In this way, it holds that $\varphi^{\prime}_{1}(u)=\varphi^{\prime}_{2}(u)$ for $u\in[\varphi^{-1}_{1}(t^{-}_{j}),\varphi^{-1}_{1}(t^{+}_{j})]$, which, combined with arguments similar to those leading to (A.23), imply

### v2

@@ Appendix A Proofs
$\zeta(z^{\prime}_{1},z^{\prime}_{2})=1/g^{\prime}(z^{\prime}_{2})>1/g^{\prime}(z_{2})=\zeta(z_{1},z_{2}),$ (A.16)
which means that any $(z_{1},z_{2})\in\mathcal{N}$ such that $z_{2}\in(a,x_{0}]$ satisfies $(z_{1},z_{2})\notin\mathcal{M}_{\zeta}$. Hence, (1) holds true. The claims (2)-(6) can be proved by similar arguments combined with the definition of $x_{1}$ and $x_{2}$. We hence omit their proofs. By the above claims (1)-(6), we know that $\mathcal{M}_{\zeta}\subseteq\cup_{i=1}^{4}\mathcal{R}_{i}$, where
@@
$\mathcal{R}_{1}=\begin{cases}\{(\tilde{z}_{1},\tilde{z}_{2})\},\,\text{ if }\beta\leq\omega_{1}(x_{2}),\\ \emptyset,\text{ if else}.\end{cases}$
$\cup_{i=2}^{4}\mathcal{R}_{i}=\begin{cases}\{(\bar{z}_{1},\bar{z}_{2})\},\,\text{ if }\beta<\omega_{2}(x_{1}),\\ \{(\hat{z}_{1}:=(g^{\prime})^{-1}_{1}(g^{\prime}(\omega_{2}^{-1}(\beta))),\hat{z}_{2}:=\omega_{2}^{-1}(\beta))\},\,\text{ if }\beta>\omega_{2}(x_{1}),\\ \{(\bar{z}_{1},\bar{z}_{2}),(\hat{z}_{1},\hat{z}_{2})\},\,\text{ if }\beta=\omega_{2}(x_{1}).\end{cases}$
Therefore, we have the following observations:
  • If $g^{\prime}(\omega_{1}^{-1}(\beta))<g^{\prime}(\omega_{2}^{-1}(\beta))$ and $\beta<\omega_{2}(x_{1})$, then $\emptyset\neq\mathcal{M}_{\zeta}\subseteq\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2})\}$, and
@@
  • If $g^{\prime}(\omega_{1}^{-1}(\beta))>g^{\prime}(\omega_{2}^{-1}(\beta))$ and $\beta<\omega_{2}(x_{1})$, then $\emptyset\neq\mathcal{M}_{\zeta}\subseteq\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2})\}$, and
@@
  • If $g^{\prime}(\omega_{1}^{-1}(\beta))=g^{\prime}(\omega_{2}^{-1}(\beta))$ and $\beta<\omega_{2}(x_{1})$, then $\emptyset\neq\mathcal{M}_{\zeta}\subseteq\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2})\}$, and
@@
  • $g^{\prime}(\omega_{1}^{-1}(\beta))>g^{\prime}(\omega_{2}^{-1}(\beta))$ and $\beta=\omega_{2}(x_{1})$ can not hold true simultaneously.
  • If $g^{\prime}(\omega_{1}^{-1}(\beta))=g^{\prime}(\omega_{2}^{-1}(\beta))$ and $\beta=\omega_{2}(x_{1})$, then $x_{2}\leq x_{1}$, $\emptyset\neq\mathcal{M}_{\zeta}\subseteq\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2}),(\hat{z}_{1},\hat{z}_{2})\}$.
    • $\bullet$
      If $x_{2}<x_{1}$, then $\hat{z}_{1}=\tilde{z}_{1}$ and $\hat{z}_{2}=\tilde{z}_{2}$ (since $\omega_{1}\equiv\omega_{2}$ on $x_{1},\infty)$), and
@@
    • $\bullet$
      If $x_{2}=x_{1}$, then $\beta=\omega_{1}(x_{2})$ and $(\tilde{z}_{1},\tilde{z}_{2})$ is understood as a two-point set $\{(\tilde{z}_{1},(g^{\prime})_{2}^{-1}(g^{\prime}(x_{2}))),(\tilde{z}_{1},x_{2})\}$ with $\tilde{z}_{1}=(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{1}^{-1}(\beta)))=(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta)))=\hat{z}_{1}.$ In addition,
      $\zeta(\bar{z}_{1},\bar{z}_{2})=1/g^{\prime}(x_{2})=\zeta(\tilde{z}_{1},(g^{\prime})_{2}^{-1}(g^{\prime}(x_{2})))=\zeta(\tilde{z}_{1},x_{2}).$
      Hence, $\mathcal{M}_{\zeta}=\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2})\}=\{(\tilde{z}_{1},(g^{\prime})_{2}^{-1}(g^{\prime}(x_{2}))),(\tilde{z}_{1},x_{2}),(\bar{z}_{1},\bar{z}_{2})\}.$
  • If $g^{\prime}(\omega_{1}^{-1}(\beta))<g^{\prime}(\omega_{2}^{-1}(\beta))$ and $\beta=\omega_{2}(x_{1})$, then $x_{1}<x_{2}$, $\emptyset\neq\mathcal{M}_{\zeta}\subseteq\{(\tilde{z}_{1},\tilde{z}_{2}),(\bar{z}_{1},\bar{z}_{2}),(\hat{z}_{1},\hat{z}_{2})\}$, and
@@
  • If either of the following cases is the case
@@
      $\beta>\omega_{2}(x_{1})$, $x_{1}<x_{2}$, and $\hat{z}_{2}\geq x_{2}$,
    then $g^{\prime}(\omega_{1}^{-1}(\beta))=g^{\prime}(\omega_{2}^{-1}(\beta))$ (since $\omega_{1}\equiv\omega_{2}$ on $x_{1}\vee x_{2},\infty)$), and then
    $\hat{z}_{1}=\tilde{z}_{1},\hat{z}_{2}=\tilde{z}_{2},$
    i.e., $\cup_{i=1}^{4}\mathcal{R}_{i}=\{(\tilde{z}_{1},\tilde{z}_{2})\}$. Hence, $\mathcal{M}_{\zeta}=\{(\tilde{z}_{1},\tilde{z}_{2})\}$.
  • If $\beta>\omega_{2}(x_{1})$, $x_{1}<x_{2}$, and $\hat{z}_{2}<x_{2}$, then
@@
$=[g^{-}(0)-g^{+}(0)](\theta_{1}^{+})^{2}>0.$ (A.20)
It is seen that the function $g^{\prime\prime}(x)$ is strictly increasing on $(a,\infty)$, which together with (A.6) implies that $g^{\prime\prime}(x)>0$ on $(a,\infty)$. The proof is complete.
@@
      If $a>a_{1}$, one has $g^{\prime\prime}(x)<0$ on $0,a_{1})$ and $g^{\prime\prime}(x)>0$ on $(a_{1},a]$.
@@
    $=[g^{-}(0)-g^{+}(0)](\theta_{1}^{+})^{2}>0,$ (A.21)
    where we have used the facts that $\theta_{1}^{+}\leq\theta_{2}^{+}$ and $g^{-}(0)>g^{+}(0)>0$. It is seen that the function $g^{\prime\prime}(x)$ is strictly increasing on $(a,\infty)$, which together with ((2)) implies that $g^{\prime\prime}(x)>0$ on $(a,\infty)$.
@@
We exclusively present the proof for case (iv) of Subsection 3.1, as proofs for the remaining cases of Subsections 3.1-3.4 adhere to patterns that are either similar or simpler in nature. Recall that $\omega_{1}$ coincides with $\omega_{2}$ on $x_{1}\vee x_{2},\infty)$. By Theorem 3.1, one has
@@
    $z_{2}-z_{1}=\begin{cases}\omega_{2}^{-1}(\beta)-(g^{\prime})_{3}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),&\beta\in[\omega_{1}(x_{2}),\omega_{2}(x_{1})),\\ \omega_{2}^{-1}(\beta)-(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),&\beta\in[\omega_{2}(x_{1}),\infty),\\ \omega_{1}^{-1}(\beta)-(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{1}^{-1}(\beta))),&\beta\in[0,\omega_{1}((g^{\prime})_{2}^{-1}(g^{\prime}(x_{0})))),\end{cases}$ (A.23)
    which implies that $z_{2}-z_{1}$ is increasing in $\beta$ on $0,\omega_{1}((g^{\prime})_{2}^{-1}(g^{\prime}(x_{0}))))\cup[\omega_{2}(x_{2}),\infty)$.
@@
    $z_{2}-z_{1}=\omega_{1}^{-1}(\beta)-(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{1}^{-1}(\beta))),\quad\beta\in[0,\omega_{1}((g^{\prime})_{2}^{-1}(g^{\prime}(x_{0}))))\cup[\omega_{1}(x_{1}),\infty),$ (A.24)
    which implies that $z_{2}-z_{1}$ is increasing in $\beta$ on $0,\omega_{1}((g^{\prime})_{2}^{-1}(g^{\prime}(x_{0}))))\cup[\omega_{1}(x_{1}),\infty)$.
@@
    $z_{2}-z_{1}=\begin{cases}\omega_{2}^{-1}(\beta)-(g^{\prime})_{3}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),&\beta\in[0,\omega_{2}(a_{5})]\cup[\omega_{2}(x_{2}),\omega_{2}(x_{1})),\\ \omega_{2}^{-1}(\beta)-(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),&\beta\in[\omega_{2}(x_{1}),\infty),\end{cases}$ (A.25)
    which implies that $z_{2}-z_{1}$ is increasing in $\beta$ on $0,\omega_{2}(a_{5})\cup[\omega_{2}(x_{2}),\infty)$.
@@
    $z_{2}-z_{1}=\begin{cases}\omega_{2}^{-1}(\beta)-(g^{\prime})_{3}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),&\beta\in[0,\omega_{2}(a_{5})],\\ \omega_{1}^{-1}(\beta)-(g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{1}^{-1}(\beta))),&\beta\in[\omega_{2}(x_{1}),\infty),\end{cases}$ (A.26)
    which implies that $z_{2}-z_{1}$ is increasing in $\beta$ on $0,\omega_{2}(a_{5})]\cup[\omega_{2}(x_{1}),\infty)$.
In the sequel, we only check the increasing property of $\max_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ and $\min_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ (with respect to $\beta$) over the interval $\omega_{1}((g^{\prime})_{2}^{-1}(g^{\prime}(x_{0}))),\omega_{1}(x_{2})]$ under the above Case (a), because the corresponding proofs needed for the above Cases (b)-(d) are quite similar.
@@
Hence, it holds that $\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),\omega_{1}(x_{2})]=\Lambda^{+}\cup\Lambda^{-}\cup\Lambda^{0}.$ We next demonstrate that $\Lambda^{\pm}$ can each be expressed as a countable union of disjoint open intervals, and, $\Lambda^{0}$ can be expressed as a countable union of disjoint closed intervals. For any $\beta_{0}\in\Lambda^{+}$, we have $g^{\prime}(\omega^{-1}_{1}(\beta_{0}))>g^{\prime}(\omega^{-1}_{2}(\beta_{0}))$. According to the sign-preserving property of continuous functions, there exists a constant $\epsilon>0$ such that $(\beta_{0}-\epsilon,\beta_{0}+\epsilon)\subseteq[\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),\omega_{2}(x_{2})]$ and for any $\beta\in(\beta_{0}-\epsilon,\beta_{0}+\epsilon)$, it holds that $g^{\prime}(\omega^{-1}_{1}(\beta))>g^{\prime}(\omega^{-1}_{2}(\beta))$. Denote
@@
$\hskip 14.22636pt\text{and, }g^{\prime}(\omega^{-1}_{1}(\beta))>g^{\prime}(\omega^{-1}_{2}(\beta))\text{ for any }\beta\in(\beta_{0}-\epsilon,\beta_{0}+\overline{\epsilon})\},$ (A.28)
@@
$\hskip 14.22636pt\text{and, }g^{\prime}(\omega^{-1}_{1}(\beta))>g^{\prime}(\omega^{-1}_{2}(\beta))\text{ for any }\beta\in(\beta_{0}-\underline{\epsilon},\beta_{0}+\epsilon)\}.$ (A.29)
With the continuity of $g^{\prime}$ on $\mathbb{R}_{+}$ and $(\omega^{-1}_{i})_{i=1,2}$ on $\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),\omega_{1}(x_{2})]$, we derive that $g^{\prime}(\omega^{-1}_{1}(\beta_{0}-\underline{\epsilon}(\beta_{0})))=g^{\prime}(\omega^{-1}_{2}(\beta_{0}-\underline{\epsilon}(\beta_{0})))$ and $g^{\prime}(\omega^{-1}_{1}(\beta_{0}+\overline{\epsilon}(\beta_{0})))=g^{\prime}(\omega^{-1}_{2}(\beta_{0}+\overline{\epsilon}(\beta_{0})))$. Additionally, it holds that $\beta_{0}-\underline{\epsilon}(\beta_{0})\leq\beta_{0}-\epsilon<\beta_{0}<\beta_{0}+\epsilon\leq\beta_{0}+\overline{\epsilon}(\beta_{0}),$ and $(\beta_{0}-\underline{\epsilon}(\beta_{0}),\beta_{0}+\overline{\epsilon}(\beta_{0}))$ is a non-empty subset of $\Lambda^{+}$. In this vein, for any $\beta_{0}\in\Lambda^{+}$, one can find a non-empty open subset of $\Lambda^{+}$ that contains $\beta_{0}$. In addition, according to the denseness of rational numbers in the set of real numbers, there is at least one rational number in $(\beta_{0}-\underline{\epsilon}(\beta_{0}),\beta_{0}+\overline{\epsilon}(\beta_{0}))$. However, rational numbers are countable, implying that $\Lambda^{+}$ is a countable union of disjoint open intervals. Using similar arguments, one can obtain that $\Lambda^{-}$ is also a countable union of disjoint open intervals. Furthermore, it follows from $\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),\omega_{1}(x_{2})]=\Lambda^{+}\cup\Lambda^{-}\cup\Lambda^{0}$ that $\Lambda^{0}$ is a countable union of disjoint closed intervals. More specifically, the three sets $\Lambda^{+}$, $\Lambda^{-}$ and $\Lambda^{0}$ can be expressed as
@@
We next demonstrate that $\max_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ and $\min_{(z_{1},z_{2})\in\mathcal{M}_{\zeta}}(z_{2}-z_{1})$ are indeed increasing functions of $\beta$ on $\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0}))),{\omega_{1}}(x_{2})]$ under Case (a). Actually, by definition, one knows that if $\beta=\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0})))$, then $g^{\prime}(\omega^{-1}_{1}(\beta))=g^{\prime}(x_{0})<g^{\prime}(\omega^{-1}_{2}(\beta))$, which yields $\omega_{1}((g^{\prime})^{-1}_{2}(g^{\prime}(x_{0})))\in\Lambda^{-}$, implying that $\Lambda^{-}\neq\emptyset$. Similarly, if $\beta=\omega_{1}(x_{2})$, then $g^{\prime}(\omega^{-1}_{1}(\beta))=g^{\prime}(x_{2})>g^{\prime}(\omega^{-1}_{2}(\beta))$, which yields $\omega_{1}(x_{2})\in\Lambda^{+}$, implying that $\Lambda^{+}\neq\emptyset.$ Hence, one has $\Lambda^{0}\neq\emptyset.$ Then by (A.27), a point $(z_{1},z_{2})\in\mathcal{M}_{\zeta}$ satisfies that
@@
It can be checked that the functions $\varphi_{1}$ and $\varphi_{2}$ are strictly increasing on $g^{\prime}(x_{0}),g^{\prime}(x_{2})]$, and inherit differentiability from $\omega_{1}$, $\omega_{2}$, $(g^{\prime})_{2}^{-1}$ and $(g^{\prime})_{4}^{-1}$. In addition, it holds that
@@
[…]
@@
[…]